A continuous random variable is uniformly distributed between 20 and 120. Find the following probabilities –...

  1. A continuous random variable is uniformly distributed between 20 and 120.
  1. Find the following probabilities –
  1. P(X<70)
  2. P(X>50)
  3. P(X=50)
  4. P(30<X<90)
  1. What are the mean and standard deviation of this distribution?
Answer #1

TOPIC:Uniform distribution.

- Define, xr Onif (20, 120) So, the palf of xis- (2) { 120-20 16:20 X$120. ( 0 ; otherwise. ; otherwise. af we need O P(X< 79③ we needs XP ( x 750). TX 750 120 dx 50 a 20 dx . 100 = to do. [2] 120 50 120-50 100 - 120-5 10.40 i we need, P ( x = 50). IWe need =) P (30 < x <90). Jxdx J 30 30 - Too [x700 go-30 100 210.60 / (An). by we need, mean (x). = E(X). 120 x Jx) dx. = 1Again, E (x2). 1120 x² jeni dr. 20 2.0 too [2 = 5733. 3333 i var (x). = E(X2) - E²(X). = 833. 3333. standard deviation (x). =

(Rounded to four decimal places).